How can I find an equation of the line whose graph includes the following two points. (-1,4), (-3,-3).
I think I have an idea on how to approach it, but I am uncertain.

Answer by rfer(16322)

slope
m=(-3-4)/(-3--1)
m=-7/-2
m=3.5
-------------
y=mx+b
-3=3.5*-3+b
-3=-10.5
7.5=b
----------------
y=3.5x+7.5